OpenStudy (anonymous):

find the inverse of (f)x= 7x - 3. what is the value of the composition of the function and its inverse

OpenStudy (jdoe0001):

hmmm did you find the inverse?

OpenStudy (anonymous):

no idk how

OpenStudy (jdoe0001):

\(\bf f(x)= {\color{red}{ y}}=7{\color{blue}{ x}} - 3\qquad inverse\implies {\color{blue}{ x}}=7{\color{red}{ y}} - 3\) to find the inverse relation of the function, as you can see, we simply switch about the variables then solve for "y", what does that give you?

OpenStudy (anonymous):

would it be (x+3)/7= y ?

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

thank you!

OpenStudy (jdoe0001):

\(\bf f(x)= {\color{red}{ y}}=7{\color{blue}{ x}} - 3\qquad inverse\implies {\color{blue}{ x}}=7{\color{red}{ y}} - 3\\\quad\\\implies \cfrac{x+3}{7}=y=f^{-1}(x)\) so now "what is the value of the composition of the function and its inverse" ?

OpenStudy (anonymous):

I have no idea

OpenStudy (jdoe0001):

one sec

OpenStudy (jdoe0001):

so the composition of f(x) and it's inverse function will be \(\bf f(x)= {\color{red}{ y}}=7{\color{blue}{ x}} - 3\qquad inverse\implies {\color{blue}{ x}}=7{\color{red}{ y}} - 3 \\ \quad \\ \implies {\color{green}{ \cfrac{x+3}{7}}}=y=f^{-1}(x) \\ \quad \\ f(\quad f^{-1}(x)\quad )=7x - 3\implies f\left({\color{green}{ \cfrac{x+3}{7}}}\right)= \cancel{7}\left({\color{green}{ \cfrac{x+3}{\cancel{7}}}}\right)-3 \\ \quad \\ f\left({\color{green}{ \cfrac{x+3}{7}}}\right)=x\cancel{+3-3}\)

OpenStudy (jdoe0001):

keep in mind that \(\bf \textit{composition of function and inverse is always }=x \\ \quad \\ \textit{that is}\qquad f(f^{-1}({\color{blue}{ x}}))={\color{blue}{ x}}\)

OpenStudy (anonymous):

is that the answer?

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

thank you so much!

OpenStudy (jdoe0001):

yw